OpenStudy (anonymous):

A jumper reaches 110 cm what is the final velocity? The equation is asked after the discussion of free fall so I am assuming the acceleration is 9.8m/s^2 Please Help!!!!!!

OpenStudy (turingtest):

the kinematic to remember here is\[v_f^2=v_i^2+2ad\]in the case of gravity being the accelerating agent we have\[v_f^2=v_i^2-2gh\]I don't really know what final velocity you want though. at the top of his jump his velocity is 0m/s (because he is not moving at the moment he is at the top of his jump) when he gets to the ground it will be back to vi, so\[0=v_i^2-2gh\]which you can solve for vi=vf